# Detectability of dark matter subhalo impacts in Milky Way stellar streams

### v2

@@ 2.3 Simulated impacts
To generate the impacted stream data, we compute the 6 observables at the central value of $\psi$ for each bin using the analytic model. Then for each value in each bin, we add a randomly sampled Gaussian error based on Eq. 5 and 6. For simplicity, we assume the number of stars per bin, $n$, to be the average value across bins and thus a constant. For $n\gg 1$, the error introduced by this approximation is negligible. The right panel of Fig. 1 shows a comparison between the observables and errors generated from a simulated perturbed stream (blue) versus the observables from analytic model and errors from the dispersion of smooth stream simulation (orange). The simulated perturbed streams are obtained using gala by injecting the subhalo potential for a short period of time ($\pm$ 20 Myr) around the time of impact. We can see that the errors from smooth stream simulation and perturbed stream simulation are very similar. Also, the differences in the observables between the simulation and analytic model are comparable to the error bars. This validates our method to rapidly generate observables for a perturbed stream from the analytic model. Note also that we have not included any observational error in this figure; including such errors will render some observables more useful than others as we will see later.
@@ 3 Observational scenarios
We first consider the distribution of magnitudes of stars in a typical cluster, which in turn depends on the mass distribution. We assume that the stars in the cluster are distributed according to the Chabrier power-law initial mass function (IMF) (Chabrier, 2001). Then the mass distribution is further convolved with a MIST isochrone (Dotter, 2016; Choi et al., 2016) with age = 10 Gyr and [Fe/H] $=-2$, corresponding to old, metal-poor stars. This gives the distribution of the magnitudes of stars in various wavelength filters. We implement this using the package imf [^2] https://github.com/keflavich/imf for the initial mass function and package minimint (Koposov, 2023) for the MIST isochrone.
Given the distribution of magnitudes, we can generalize Eqs. 5-6, which only account for internal stream dispersion, to include measurement uncertainty and survey limits. As discussed below in Sec. 3.1, for stream density ratio $\rho/\rho_{0}$ and position $\overline{\Delta\theta}$, observational errors can be neglected, and the effect can simply be accounted for by the fraction of observable stars, $n_{\rm obs}/n$. For the other kinematic quantities $\overline{X}$, we define an effective per-star uncertainty $\sigma_{X}^{\star}$:
$\sigma_{\overline{X}}^{2}$ $\equiv\frac{\sigma_{X}^{\star 2}}{n}.$ (14)
The quantity $\sigma_{X}^{\star 2}$ folds in multiple effects: internal dispersion of the stream, the fraction of observable stars, and measurement uncertainties in the observations. These combine in a non-linear way to give the total error, so we have found it useful to capture all effects in a single quantity. Here $n$ is the total number of stars per bin, given by
$n=\lambda\cdot(\psi~{}\mathrm{bin~{}size}),$ (15)
where $\lambda$ is the average stellar density. This is the number of stars per bin in the limit of perfect observations and according to Eq. 12, only depends on total stellar mass $M_{\mathrm{stellar}}$ and stream length $l$. Thus, we refer to $\sigma_{X}^{\star 2}$ as an effective uncertainty per star, if all stars of the stream were measured. Note that for simplicity, we assume $\lambda$ and $n$ to be constant across bins.
We now discuss how the observational errors are incorporated into each observable, as well as the dependence of the uncertainties $\sigma_{X}^{\star 2}$ on stream width $\sigma_{\theta}$ and distance to the Galactic Center $r_{0}$. Throughout this section, we keep the age and metallicity fixed as these are higher order effects in estimating detectability.
@@ 3.1.1 Number density
Here $n_{\mathrm{obs}}$ is the number of observable stars per bin below a certain magnitude cutoff, while $n$ is the number of actual stars per bin, observable or not. We have written the error in this way to split it into two factors. The first factor, $1/n$, is the error in the limit of perfect observations. The second factor, $n/n_{\mathrm{obs}}$, only depends on the observational scenario, or more explicitly, the magnitude cutoff we use. Since observational errors on the angular positions of stars are typically much smaller than our default bin size of 2 degrees, we assume that stars can be perfectly assigned to a bin.
The magnitude cutoff can be made based on (1) kinematic (proper motions) identification of stream stars, (2) spectroscopic identification or (3) photometric measurements only. In the first case, we consider kinematic information from the Gaia data  (Gaia Collaboration et al., 2021) (with magnitude $G<20.7$). In the second case, we consider spectroscopy from the Via project (with magnitude $G<24$). We assume these provide high efficiency of stream star selection with negligible contamination of non-stream stars.
@@ 3.1.2 Positions
In considering the error for the observable $\overline{\Delta\theta}$, we use the same magnitude cutoff as the density observable in the previous section. So the percentage of observable stars $n_{\mathrm{obs}}/n$ in Fig. 3 applies here as well. Among the observable stars, the observational errors for star positions on the sky are negligible compared to the stream width. For instance, the position uncertainty with LSST is $<$ 0.1 arcsec for observations down to an apparent r-band magnitude of $r=24$. This contribution to the error, $\sigma_{\theta,\mathrm{obs}}=0.1~{}\mathrm{arcsec}=2.78~{}\times 10^{-5}~{}\text{deg}$, is far less than the stream width, $\sigma_{\theta,\text{disp}}\sim 0.1-2$ deg. We will therefore only account for the reduction in the number of stars from the magnitude cutoff, and not consider the observational errors on those stars. The error for $\overline{\Delta\theta}$ can then be written as:
@@
Figure 4: Top left: Observational errors on radial velocity as a function of the apparent magnitude G for the Via survey (green) and the DESI survey (orange) and as a function of apparent magnitude V for 4MOST (purple). Top right: The cumulative apparent magnitude distribution for Via (green), DESI (orange) and 4MOST (purple) at different stream distances $r_{0}$. Bottom left: The effective per-star uncertainty $\sigma_{v_{r}}^{\star 2}$ as defined in Eq. 21 as a function of magnitude cutoff for Via (green), DESI(orange) and 4MOST (purple) at distance of $r_{0}=10$ kpc. The error for different widths of streams are shown in different levels of darkness. Bottom right: The same thing as bottom left but at distance of $r_{0}=30$ kpc.
@@ 3.1.3 Radial velocity
For the velocity observables, the observational errors are no longer negligible compared to the internal dispersion. Furthermore, compared to the statistical error arising solely from stream dispersion, a difference in this case is that observational errors vary significantly from star to star depending on stellar properties, primarily the magnitude. Rather than using a uniformly weighted average $\overline{X}=\sum_{i}X_{i}/n$ to estimate the mean, we will use an inverse-variance weighted estimator when taking both dispersion and observational errors into account:
@@
$\sigma^{2}_{\overline{X}}$ $=\frac{1}{n}\frac{1}{\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{bin}}}$ (19)
@@
$\approx\frac{1}{n}\frac{1}{\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{cluster}}}.$ (20)
In the second line, we replaced the average over stars in a given bin with an average over all stars in the cluster. This therefore neglects fluctuations in the variance across bins and is a reasonable approximation for $n\gg 1$.
As before, we have split the error into two parts. The first part $1/n$ only depends on stream properties. The second part depends on the internal dispersion and the distribution of observational errors. The quantity $\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{cluster}}$ is averaged over all stars of the cluster by assuming $\sigma_{X_{i},\mathrm{obs}}=\infty$ for unobservable stars (stars with magnitudes above the cutoff) thereby encoding the information about the magnitude cutoff. The effective uncertainty as defined in Eq. 14 is then given by,
@@
$\sigma_{X}^{\star 2}\equiv\frac{1}{\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{cluster}}}.$ (21)
@@
Figure 5: Top left: Observational errors on $v_{z}$ as a function of the apparent magnitude G for Gaia (green) and as a function of the apparent magnitude r for LSST (blue) for different stream distances $r_{0}$. Top right: The cumulative apparent magnitude distribution for Gaia and LSST at different stream distances $r_{0}$. Bottom left: The effective per-star uncertainty $\sigma_{v_{z}}^{\star 2}$ as defined in Eq. 21 as a function of magnitude cutoff for Gaia (red) and Gaia + LSST (blue), at distance of $r_{0}=10$ kpc. The errors for different widths of streams are shown in different levels of darkness. Bottom right: Same as bottom left but at distance of $r_{0}=30$ kpc.
@@
Now we discuss how to calculate this for $X=v_{r}$. Radial velocities are assumed to be measured by one of the 4MOST survey (de Jong et al., 2012), the DESI survey (Cooper et al., 2023), or the Via project. The observational errors on individual stars as a function of apparent magnitude is shown in the top left panel of Fig. 4 for each survey. These also depend on the metallicity of the stars, and can change up to a factor of 2 with a change in metallicity [Fe/H] of 1. For 4MOST (de Jong et al., 2012) and Via (Chandra, 2024), we use the error estimate for the metallicity of [Fe/H]=$-2$. For DESI (Koposov et al., 2024), we download the dataset from DESI Data website [^3] https://data.desi.lbl.gov/doc/releases/edr/vac/mws/, and extract the mean errors from stars with metallicity [Fe/H] between -2.5 and -1.5. In the same panel, we also plot the internal dispersion in radial velocity for different widths of streams as gray dashed lines. This can be obtained from the simulated streams introduced in Sec. 2.2.
@@
Given the stream dispersion, observational errors at each magnitude (upper left panel of Fig. 4), and the distribution of magnitudes (upper right panel of Fig. 4), we can calculate the effective per-star uncertainty in Eq. 21 for different surveys. The bottom panels of Fig. 4 show the resulting $\sigma_{v_{r}}^{\star 2}$ as a function of the magnitude cutoff for 4MOST (purple), DESI (orange) and Via (green).
For each survey, to illustrate the dependence on stream width $\sigma_{\theta}$, we show three lines representing $\sigma_{\theta}=0.1^{\circ}$, $\sigma_{\theta}=0.3^{\circ}$ and $\sigma_{\theta}=1^{\circ}$ respectively. To show the variation with distance to the stream $r_{0}$, we split the bottom panel into two subplots, $r_{0}=10$ kpc (bottom left) and $r_{0}=30$ kpc (bottom right). We can see that in all cases, the errors reach a plateau before the maximum cutoff magnitude for 4MOST, DESI or Via. This indicates that it is possible to achieve similar performance even if these surveys do not measure all stars in the stream down to the maximum magnitude cutoff. The faint end of the stellar distribution is only marginally helpful in detecting subhalos. Nevertheless, we will still use the maximum possible cutoff magnitude to minimize the errors.
@@ 3.1.4 Proper motions
Figure 6: The effective per-star uncertainty $\sigma_{X}^{\star 2}$, as defined in Eq. 21, for $X=v_{r}$ with Via, 4MOST, and DESI, as well as $X=v_{z}$ with Gaia and LSST. The uncertainty is shown as a function of stream distance $r_{0}$ and stream width $\sigma_{\theta}$, where the stream width also determines velocity dispersion $\sigma_{\theta}\propto\sigma_{v_{r}}\propto\sigma_{v_{z}}$.
For proper motions, the observational errors are also not negligible compared to internal dispersion, and we follow the same framework as for the radial velocity to compute the effective per-star uncertainty in Eq. 14.
@@
With information from both of the top panels of Fig. 5, we calculate the effective per-star uncertainty in Eq. 21 for $X=v_{z}$ under assumptions of different magnitude cutoffs. In the bottom panels of Fig. 5, we show $\sigma_{v_{z}}^{\star 2}$ as a function of magnitude cutoff for Gaia (red) and Gaia + LSST (blue). In the scenario labeled “Gaia + LSST”, we use Gaia for stars with magnitude $G<20.7$ and use LSST for stars with magnitude $G>20.7$, since the observational errors for Gaia are smaller than LSST at magnitudes $G<20.7$, as can be seen in the top left panel. We again show three lines for each case, corresponding to different choices for internal dispersion, which are parameterized by stream widths ($\sigma_{\theta}=0.1^{\circ}$, $\sigma_{\theta}=0.3^{\circ}$, $\sigma_{\theta}=1^{\circ}$). Again we see that in all cases, the errors start to flatten before the maximum magnitude cutoff $G<20.7$, with the additional LSST data only giving mild improvement in increasing detectability. Nevertheless, we will still use the cutoff magnitude determined by each survey limit to maximize the number of observable stars per bin.

### v3

@@ 2.3 Simulated impacts
To generate the impacted stream data, we compute the 6 observables at the central value of $\psi$ for each bin using the analytic model. Then for each value in each bin, we add a randomly sampled Gaussian noise based on Eq. 5 and 6. For simplicity, we assume the number of stars per bin, $n$, to be the average value across bins and thus a constant. For $n\gg 1$, the error introduced by this approximation is negligible. The right panel of Fig. 1 shows a comparison between the observables and uncertainties from a simulated perturbed stream (blue) versus the observables from analytic model plus random Gaussian noises generated based on the dispersion of smooth stream simulation (orange). The simulated perturbed streams are obtained using gala by injecting the subhalo potential for a short period of time ($\pm$ 20 Myr) around the time of impact. We can see that the uncertainties from smooth stream simulation and perturbed stream simulation are very similar. Also, the differences in the observables between the simulation and analytic model are comparable to the error bars. This validates our method to rapidly generate observables for a perturbed stream from the analytic model. Note also that we have not included any observational error in this figure; including such errors will render some observables more useful than others as we will see later.
@@ 3 Observational scenarios
We first consider the distribution of magnitudes of stars in a typical cluster, which in turn depends on the mass distribution. We assume that the stars in the cluster are distributed according to the Chabrier power-law initial mass function (IMF) (Chabrier, 2001). Then the mass distribution is further convolved with a MIST isochrone (Dotter, 2016; Choi et al., 2016) with age = 10 Gyr and [Fe/H] =2=-2, corresponding to old, metal-poor stars. This gives the distribution of the magnitudes of stars in various wavelength filters. We implement this using the package imf [^2] https://github.com/keflavich/imf for the initial mass function and package minimint (Koposov, 2023) for the MIST isochrone.
Given the distribution of magnitudes, we can generalize Eqs. 5-6, which only account for internal stream dispersion, to include measurement uncertainty and survey limits. As discussed below in Sec. 3.1, for stream density ratio $\rho/\rho_{0}$ and position $\overline{\Delta\theta}$, observational errors can be neglected, and the effect can simply be accounted for by the fraction of observable stars, $n_{\rm obs}/n$. For the other kinematic quantities $\overline{X}$, we consider both the observational errors and the fraction of observable stars.
We now discuss how the observational errors are incorporated into each observable. Throughout this section, we keep the age and metallicity fixed as these are higher order effects in estimating detectability.
@@ 3.1.1 Number density
Here $n_{\mathrm{obs}}$ is the number of observable stars per bin below a certain magnitude cutoff, while $n=\lambda\cdot(\psi~\mathrm{bin~size})$ is the number of actual stars per bin, observable or not. We have written the error in this way to split it into two factors. The first factor, $1/n$, is the error in the limit of perfect observations. The second factor, $n/n_{\mathrm{obs}}$, only depends on the observational scenario, or more explicitly, the magnitude cutoff we use. Since observational errors on the angular positions of stars are typically much smaller than our default bin size of 2 degrees, we assume that stars can be perfectly assigned to a bin.
The magnitude cutoff can be made based on (1) kinematic (proper motions) identification of stream stars, (2) spectroscopic identification or (3) photometric measurements only. In the first case, we consider kinematic information from the Gaia data  (Gaia Collaboration et al., 2021) (with magnitude $G<20.7$). In the second case, we consider spectroscopy from the Via project (with maximum magnitude $G<24$). We assume these provide high efficiency of stream star selection with negligible contamination of non-stream stars.
@@ 3.1.2 Positions
In considering the error for the observable $\overline{\Delta\theta}$, we use the same magnitude cutoff as the density observable in the previous section. So the percentage of observable stars $n_{\mathrm{obs}}/n$ in Fig. 3 applies here as well. Among the observable stars, the observational errors for star positions on the sky are negligible compared to the stream width. For instance, the position uncertainty with LSST is << 0.1 arcsec for observations down to an apparent r-band magnitude of $r=24$. This contribution to the error, $\sigma_{\theta,\mathrm{obs}}=0.1~\mathrm{arcsec}=2.78~\times 10^{-5}~\text{deg}$, is far less than the stream width, $\sigma_{\theta,\text{disp}}\sim 0.1-2$ deg. We will therefore only account for the reduction in the number of stars from the magnitude cutoff, and not consider the observational errors on those stars. The error for $\overline{\Delta\theta}$ can then be written as:
@@
Figure 4: Top left: Observational errors on radial velocity as a function of the apparent magnitude G for the Via survey (green) and the DESI survey (orange) and as a function of apparent magnitude V for 4MOST (purple). Top right: The cumulative apparent magnitude distribution for Via (green), DESI (orange) and 4MOST (purple) at different stream distances $r_{0}$. Bottom left: The survey-effective uncertainty $\sigma_{v_{r}}^{\star}$ as defined in Eq. 21 and calculated according to Eq. 20 as a function of magnitude cutoff for Via (green), DESI(orange) and 4MOST (purple) at distance of $r_{0}=10$ kpc. The error for different widths of streams are shown in different levels of darkness. The survey-effective uncertainty $\sigma_{v_{r}}^{\star}$ is normalized to the total intrinsic number of stream stars and should not be interpreted as the uncertainty of an individual observed star. As the magnitude cutoff increases, a survey observes more stars and $\sigma_{v_{r}}^{\star}$ decreases. However, observing additional stars with very large individual uncertainties eventually provides diminishing returns, leading to the plateau at the end. Bottom right: The same thing as bottom left but at $r_{0}=30$ kpc. In this case, there is a rapid decrease in $\sigma^{\star}_{v_{r}}$ at $G\approx 18$, corresponding to the increased number of stars when the red clump is accessible (also visible in the top right figure).
@@ 3.1.3 Radial velocity
For the velocity observables $X=v_{r},v_{t},v_{z}$, the observational errors are no longer negligible compared to the internal dispersion. Furthermore, compared to the statistical error arising solely from stream dispersion, a difference in this case is that observational errors vary significantly from star to star depending on stellar properties, primarily the magnitude. Rather than using a uniformly weighted average $\overline{X}=\sum_{i}X_{i}/n$ to estimate the mean, we will use an inverse-variance weighted estimator when taking both dispersion and observational errors into account:
@@
$\sigma^{2}_{\overline{X}}$ $=\frac{1}{n_{\mathrm{obs}}}\frac{1}{\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{bin}}}$ (17)
Here the angular brackets refers to an average over observable stars in a given bin. Next, we approximate the average over stars in a given bin with an average over all observable stars in the cluster:
@@
$\sigma^{2}_{\overline{X}}$ $\approx\frac{1}{n_{\mathrm{obs}}}\frac{1}{\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{cluster}}}$ (18)
$=\frac{1}{n}\frac{n}{n_{\mathrm{obs}}}\frac{1}{\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{cluster}}}.$ (19)
This therefore neglects fluctuations in the variance across bins and is a reasonable approximation for $n\gg 1$. In the last line, we split the error into three parts. The first part, $1/n$, depends only on the total stellar density. The second part, $n/n_{\mathrm{obs}}$, is the fraction of observable stars as shown in Fig. 3, which encodes information about the magnitude cutoff and stream distance. The third part, $\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{cluster}}$, depends on the internal dispersion and the distribution of observational errors.
Now we define a survey-effective uncertainty $\sigma_{X}^{\star}$ which captures all survey and measurement effects, while being independent of the total stellar density (observable or not) of the stream:
@@
$\sigma_{X}^{\star 2}\equiv\frac{n}{n_{\mathrm{obs}}}\frac{1}{\langle 1/(\sigma_{X,\mathrm{disp}}^{2}+\sigma_{X,\mathrm{obs}}^{2})\rangle_{\mathrm{cluster}}}.$ (20)
The quantity $\sigma_{X}^{\star 2}$ folds in multiple effects: the fraction of observable stars, internal dispersion of the stream, and measurement uncertainties in the observations. These combine in a non-linear way to give the total error, so we have found it useful to capture all effects in a single quantity. This way, the dependence on the survey properties, including magnitude cut, is entirely captured in $\sigma_{X}^{\star}$. As the magnitude cut is increased, there are more observable stars, leading to decreasing $\sigma_{X}^{\star}$. However, this effect eventually saturates if the additional observed stars have large measurement uncertainties.
Then the total uncertainty on the mean observable is
$\sigma_{\overline{X}}^{2}\equiv\frac{\sigma_{X}^{\star 2}}{n},$ (21)
where $n$ is the total number of stars per bin, observable or not. We emphasize that $\sigma_{X}^{\star}$ is not the uncertainty of an individual observed star. Rather, the dependence of $\sigma_{X}^{\star}$ on magnitude cutoffs and distance to the stream provides a useful way to compare different surveys, independent of a stream’s stellar density.
@@
Figure 5: Top left: Observational errors on $v_{z}$ as a function of the apparent magnitude G for Gaia (red) and as a function of the apparent magnitude r for LSST (blue) for different stream distances $r_{0}$. Top right: The cumulative apparent magnitude distribution for Gaia and LSST at different stream distances $r_{0}$. Bottom left: The survey-effective uncertainty $\sigma_{v_{z}}^{\star}$ as defined in Eq. 21 and calculated according to Eq. 20 as a function of magnitude cutoff for Gaia (red) and Gaia + LSST (blue), at distance of $r_{0}=10$ kpc. The errors for different widths of streams are shown in different levels of darkness. The survey-effective uncertainty $\sigma_{v_{z}}^{\star}$ is normalized to the total intrinsic number of stream stars and should not be interpreted as the uncertainty of an individual observed star. As the magnitude cutoff increases, a survey observes more stars and $\sigma_{v_{z}}^{\star}$ decreases. However, observing additional stars with very large individual uncertainties eventually provides diminishing returns, leading to the plateau at the end. Bottom right: Same as bottom left but at $r_{0}=30$ kpc. In the bottom panels, there is a rapid decrease in $\sigma^{\star}_{v_{z}}$ at $G\approx 15$ (left) and $G\approx 18$ (right), corresponding to the increased number of stars when the red clump is accessible (also visible in the top right figure).
@@
Now we discuss how to calculate the survey-effective uncertainty for $X=v_{r}$. Radial velocities are assumed to be measured by one of the 4MOST survey (de Jong et al., 2012), the DESI survey (Cooper et al., 2023), or the Via project. The observational errors on individual stars as a function of apparent magnitude is shown in the top left panel of Fig. 4 for each survey. These also depend on the metallicity of the stars, and can change up to a factor of 2 with a change in metallicity [Fe/H] of 1. For 4MOST (de Jong et al., 2012) and Via (Chandra, 2024), we use the error estimate for the metallicity of [Fe/H]=2-2. For DESI (Koposov et al., 2024), we download the dataset from DESI Data website [^3] https://data.desi.lbl.gov/doc/releases/edr/vac/mws/, and extract the mean errors from stars with metallicity [Fe/H] between -2.5 and -1.5. In the same panel, we also plot the internal dispersion in radial velocity for different widths of streams as gray dashed lines. This can be obtained from the simulated streams introduced in Sec. 2.2.
@@
Given the stream dispersion, observational errors at each magnitude (upper left panel of Fig. 4), and the distribution of magnitudes (upper right panel of Fig. 4), we can calculate the survey-effective uncertainty in Eq. 20 for different surveys. The bottom panels of Fig. 4 show the resulting $\sigma_{v_{r}}^{\star}$ as a function of the magnitude cutoff for 4MOST (purple), DESI (orange) and Via (green).
For each survey, to illustrate the dependence on stream width $\sigma_{\theta}$, we show three lines representing $\sigma_{\theta}=0.1^{\circ}$, $\sigma_{\theta}=0.3^{\circ}$ and $\sigma_{\theta}=1^{\circ}$ respectively. To show the variation with distance to the stream $r_{0}$, we split the bottom panel into two subplots, $r_{0}=10$ kpc (bottom left) and $r_{0}=30$ kpc (bottom right). We can see that in all cases, $\sigma_{v_{r}}^{\star}$ reaches a plateau before the maximum cutoff magnitude for 4MOST, DESI or Via. This indicates that it is possible to achieve similar performance even if these surveys do not measure all stars in the stream down to the maximum magnitude cutoff. The faint end of the stellar distribution is only marginally helpful in detecting subhalos. This point is important because the ability to observe all stars down to the maximum possible cutoff is limited by observing time. For instance, Via plans to target all candidate members identified based on photometry and astrometry down to $G<21$, and is not guaranteed to observe all stars down to $G<24$. Nevertheless, we will still use the maximum possible cutoff magnitude and therefore maximum number of stars to minimize the errors.
@@ 3.1.4 Proper motions
Figure 6: The survey-effective uncertainty $\sigma_{X}^{\star}$, as defined in Eq. 20, for $X=v_{r}$ with Via, 4MOST, and DESI, as well as $X=v_{z}$ with Gaia and LSST. The uncertainty is shown as a function of stream distance $r_{0}$ and stream width $\sigma_{\theta}$, where the stream width also determines velocity dispersion $\sigma_{\theta}\propto\sigma_{v_{r}}\propto\sigma_{v_{z}}$.
For proper motions, the observational errors are also not negligible compared to internal dispersion, and we follow the same framework as for the radial velocity to compute the survey-effective uncertainty in Eq. 21.
@@
With information from both of the top panels of Fig. 5, we calculate the survey-effective uncertainty in Eq. 20 for $X=v_{z}$ under assumptions of different magnitude cutoffs. In the bottom panels of Fig. 5, we show $\sigma_{v_{z}}^{\star}$ as a function of magnitude cutoff for Gaia (red) and Gaia + LSST (blue). In the scenario labeled “Gaia + LSST”, we use Gaia for stars with magnitude $G<20.7$ and use LSST for stars with magnitude $G>20.7$, since the observational errors for Gaia are smaller than LSST at magnitudes $G<20.7$, as can be seen in the top left panel. We again show three lines for each case, corresponding to different choices for internal dispersion, which are parameterized by stream widths ($\sigma_{\theta}=0.1^{\circ}$, $\sigma_{\theta}=0.3^{\circ}$, $\sigma_{\theta}=1^{\circ}$). Again we see that in all cases, $\sigma_{v_{z}}^{\star}$ starts to flatten before the maximum magnitude cutoff $G<20.7$, with the additional LSST data only giving mild improvement in increasing detectability. Nevertheless, we will still use the cutoff magnitude determined by each survey limit to maximize the number of observable stars per bin.